karlieeeeeee:

Is (1, 1) a solution to this system of equations? 14x + y = 15 2x + 9y = 11

Convert:

Well for these try plugging the x and y coordinates into the equation: (1,1) where the first 1, is the x. And the second 1, is the y.

Convert:

So it'd look something like 14(1)+1=15 and 2(1)+9(1)=15. And if those are true, then yes, it would be a solution

Convert:

So @karlieeeeeee, can you evaluate those equations for me and tell me if they're equal?

CripQUEZZ:

Ok so (1,1) look at it and try it out first So set up he first problem like this 14(1) + (1) = 15 All you do is plug that in and you get your answer so like 14 x 1 + 1 x 1 = 15 First do (14 x 1) So it’s 14 + 1 x 1 = 15 And 1x 1 is jus 1 so then add across 14 + 1 = 15 So the first one is correct so do the same with the second one and you’ll get your answer